OpenStudy (anonymous):

Simplify: √128x^5y^2 Show work pleasee(:

OpenStudy (blurbendy):

Sqrt[128x^5y^2] = 8x^2y√2

OpenStudy (blurbendy):

oops

OpenStudy (blurbendy):

8x^2y√2x

OpenStudy (anonymous):

does it help to know that \(128=2^7\) ?

OpenStudy (anonymous):

you are taking the square root of \(2^7x^5y^2\) and you do this in your head 2 goes in to 7 3 times with a remainder of 1 so \(2^3\) comes out of the radical, one \(2\) stays in 2 goes in to 5 two times with a remainder of 1, so \(x^2\) comes out and one \(x\) stays in 2 goes in to 2 once, so \(y\) comes out you get \[2^3x^2y\sqrt{2x}\]

OpenStudy (anonymous):

There aren't 2 numbers that multiply to be 128... I don't think

OpenStudy (anonymous):

\(128=2^7\)

OpenStudy (anonymous):

it should also be pretty clear that since 128 is even, it must be two times something half of 128 is 64, so it is \(128=2\times 64\)

OpenStudy (anonymous):

Oh! Okay thank you

OpenStudy (anonymous):

@itsallycatt see if you can understand the method i wrote above, it will make your life much simpler when you do these. it is a lot of words, but that is because i was trying to write what i was actually thinking. it is not hard to understand